Respected sir,
[1] A man loses Rs.20 by selling some toys at the rate of Rs.3 per piece and gains Rs.30,if he sells them at Rs. 3.25 per piece.The number of pieces sold by him are ?
waiting for yurs reply sir..plz solve my difficulty.
Regards
Binita

Answer by RAY100(1637)

let b = base price (cost)
let n = number of pieces
.
1) b-20 =3n
.
2) b+30 = 3.25 n
.
subtract (1) from (2)
.
50 = .25 n
.
n= 200 pieces
.
subst into (1) b-20 = 3(200) = 600
.
b=620 R ,,,,,
.
checking
1) 620 - 20 = 3 * 200
600 = 600,,,,ok
2) 620 +30 = 3.25 * 200
650 = 650 ,,,, ok